ROTATIONAL SPECTRA OF THE N_{2}O-paraD_{2} COMPLEX

\underline{JEN NICOLE LANDRY} AND WOLFGANG JÄGER, Department of Chemistry, University Of Alberta, Edmonton, AB T6G 2G2, Canada.

Using a pulsed-nozzle Fourier-transform microwave spectrometer, rotational transitions of the N_{2}O-paraD_{2} complex were measured in the 5 - 21 GHz frequency region. For each of the ^{14}N^{14}NO-pD_{2}, ^{15}N^{14}NO-pD_{2}, ^{14}N^{15}NO-pD_{2} and ^{15}N^{15}NOpD_{2} isotopmers, 4 transitions were recorded. In addition, the nuclear quadrupole hyperfine structure due to the presence of two ^{14}N (I=1) nuclei and pD_{2} (I_{tot}=1) were detected and analyzed. Three potential energy surfaces with different orientations of the pD_{2} unit relative to N_{2}O were calculated at the CCSD(T) level of theory. The aug-cc-pVTZ basis set was used for all atoms in the complex. Midbond functions were used to complement the basis set. The strategy is to construct a hybrid surface that can reproduce the observed transition frequencies.